OpenStudy (unicwaan):

I need to find the domain and range, x and y intercepts, horizontal and vertical asymptotes. Can someone help me through it please?

OpenStudy (anonymous):

Hi what is your function?

OpenStudy (unicwaan):

There should be a picture attached but it's: X^2 + x - 2 divided by x^2 - 3x - 4

OpenStudy (anonymous):

first thing, you should factor the num and denom of your function\[f(x)=\frac{x^2+x-2}{x^2-3x-4}\]have you tried to do so?

OpenStudy (unicwaan):

I did, the numerator factors into (x-1)(x+2) and the deniminator factors into (x-4)(x+1)

OpenStudy (anonymous):

very good, so what will be the domain considering the denom?

OpenStudy (unicwaan):

The domain would be all reals except x=4 and x=-1?

OpenStudy (anonymous):

quite right

OpenStudy (anonymous):

let's work on range of the function

OpenStudy (unicwaan):

alright

OpenStudy (unicwaan):

Is the range all reals?

OpenStudy (anonymous):

thats right, do you have a reasoning for that?

OpenStudy (unicwaan):

Well by the graph it extends infinitely up and down

OpenStudy (unicwaan):

and because quadratic functions' range will always be all reals

OpenStudy (anonymous):

that's right, but what if we don't know about the graph

OpenStudy (anonymous):

note that quadratics don't cover all reals

OpenStudy (unicwaan):

Oh okay, then I do not a reasoning

OpenStudy (unicwaan):

do not have a reasoning*

OpenStudy (anonymous):

function is continuous on the interval (-1,4) and\[\lim_{x \to -1^{+}} f(x)=\infty \]\[\lim_{x \to 4^{-}} f(x)=-\infty \]therefore range is all of real numbers

OpenStudy (anonymous):

does that make sense?

OpenStudy (unicwaan):

So when you plugged in the ordered pair, the output was infinite basically?

OpenStudy (anonymous):

thats right

OpenStudy (unicwaan):

okay that makes sense

OpenStudy (anonymous):

for example\[\lim_{x \to -1^{+}} f(x)=\frac{1 \times -2}{-5 \times 0^{+}}=\infty\]

OpenStudy (anonymous):

ok what are x and y intercepts?

OpenStudy (unicwaan):

Alright so the y intercept = 1/2 because you plug in 0 for x to find y. The x intercept = -2,1 because they are the zeros of the numerator

OpenStudy (anonymous):

very right, thanks

OpenStudy (anonymous):

and what about asymptotes

OpenStudy (unicwaan):

The horizontal asymptotes is 1 because the degree of each function are the same and therefore, must divide the leading coeeficients. The vertical asymptotes are y= -1,4 because they are the zeros of the denominator.

OpenStudy (anonymous):

right, you can put your words in mathematical phrases

OpenStudy (anonymous):

now you can graph your function

OpenStudy (unicwaan):

Okay so for the asympotes I put dotted lines to represent them, and then I plugged in points according to get this graph Poor drawing but the basic jist of it :/